# Odd Dipole Screening in Radial Inflation

### v1

@@ IV Summary and discussion
The aim of this paper was to explore further the consequences of Chiral symmetry breaking in the mechanical response of amorphous solids. This symmetry breaking is additional to the breaking of translational symmetry. The latter is responsible for the introduction of a typical screening length, which in the notation of the present paper is $\kappa_{e}^{-1}$. Translational symmetry breaking due to dipole screening is also the mechanism behind the hexatic phase transition [19] and the Kosterlitz-Thouless transition [20]. The appearance of this Chiral symmetry breaking seems novel to the mechanical response of amorphous solids, and is related to irreversibility and energy loss (or gain) in a closed loop of straining an amorphous solids. It can be shown that the presence of the odd tensor ${\bm{\Gamma}}$ in the form of Eq. (6) is sufficient to render a closed loop of straining dissipative, in the sense that energy is either lost or gained by the material [7]. It is worthwhile to remark that a similar notion of “Odd Elasticity" was coined recently in the context of materials whose mechanics is not derived from a potential [21]. The present formulation thus may be referred to as “Odd Plasticity", and it also connects to the lack of energy conservation.
@@
The methods employed above translate naturally to experimental data. All that one needs is an accurate map of the displacement field, something that was shown to be very doable in experiments using transparent plastic disks, see for example Ref. [6]. Once the displacement field is given, the protocol followed in the previous section is relatively straightforward. It seems quite obvious that the interest in the presence of Chiral symmetry breaking warrants a serious experimental attempt to corroborate the simulation results shown above.
@@
An issue that was not explicitly discussed in the present paper is the transition from elastic response to anomalous one. It was shown before that given glass formers or granular system can show both behaviors with an interesting transition between them, as a function of a control parameter like pressure, [4]. We now expect that the transition will be accompanied by both translation and Chiral symmetry breaking, requiring additional scrutiny of the nature of this transition. This is work in progress in our group that will be reported in the near future.
@@
Finally, dynamical response to oscillatory or any form of dynamical straining are known to show dipole screening as well, cf. Ref. [5]. The presence of Chiral symmetry breaking and it consequences needs to be examined for such protocols as well.
@@ Appendix A Analytic solutions of the equations
Substituting the values of $Z$ from above equations into Eqn.(A), we obtain the values of $\kappa$ defining the transformation in Eqn.(33). Therefore, from equations (31, and 33) we have
@@
Thus Eqns. (65 and 66) reduce to
@@
Now we can write the values of $m_{1}$ and $m_{2}$ from Eqs.(62 and 64)

### v3

@@ IV Summary and discussion
It is important to distinguish the present phenomena from those described by Cosserat Elaticity Rueger et al. (2019) and Odd Elasticity Scheibner et al. (2020). Both these theories enlarge classical elasticity, that contains only shear and bulk moduli, to describe more general situations. The former theory allows the presence of internal rotational degrees of freedom; In an isotropic Cosserat solid there are six elastic constants rather than two. The latter theory allows active constituents that can propel themselves, resulting in a non-symmetric elastic tensor. Our discussion pertains to plastic responses that are accompanied by emergent topological charges in the response, rather than in the underlying structure of the material. We find additional moduli, but they do not appear in the relation between stress and strain, but rather in the relation between quadrupolar and dipolar fields and the displacement field.
A central aim of this paper was to explore further the consequences of Chiral symmetry breaking in the mechanical response of amorphous solids. This stems from the constitutive relation between the dipole and the displacement that the theory predicts, i.e.
${\bm{{\mathcal{P}}}}=-{\bm{\Gamma}}{\bm{d}}\ .$ (27)
@@
Observing the explicit form Eq. (9),we see that for $\kappa_{o}=0$ the prediction is that the dipole field should be co-linear with the displacement field. With $\kappa_{o}\neq 0$ we expect rotation, with an angle between the inverse displacement field and the dipole field as seen in Eq. 7. This angle can be positive or negative, but given for any particular displacement field in a given system. This symmetry breaking is additional to the breaking of translational symmetry. The latter is responsible for the introduction of a typical screening length, which in the notation of the present paper is $\kappa_{e}^{-1}$. Translational symmetry breaking due to dipole screening is also the mechanism behind the hexatic phase transition Halperin and Nelson (1978) and the Kosterlitz-Thouless transition Kosterlitz (2016). The appearance of this Chiral symmetry breaking seems novel to the mechanical response of amorphous solids, and it was argued in Ref. Cohen et al. (2024) that it is related to irreversibility and energy loss (or gain) in a closed loop of straining an amorphous solids. It is worthwhile to remark that the similar notion of “Odd Elasticity" that was discussed above was been coined in the context of active materials whose mechanics is not derived from a potential Braverman et al. (2021). The present formulation thus may be referred to as “Odd Plasticity", as it also connects to the lack of energy conservation.
@@
The methods employed above translate naturally to experimental data. All that one needs is an accurate map of the displacement field, something that was shown to be very doable in experiments using transparent plastic disks, see for example Ref. Mondal et al. (2022). Once the displacement field is given, the protocol followed in the previous section is relatively straightforward. It seems quite obvious that the interest in the presence of Chiral symmetry breaking warrants a serious experimental attempt to corroborate the simulation results shown above.
@@
An issue that was not explicitly discussed in the present paper is the transition from elastic response to anomalous one. It was shown before that given glass formers or granular system can show both behaviors with an interesting transition between them, as a function of a control parameter like pressure, Jin et al. (2024). We now expect that the transition will be accompanied by both translation and Chiral symmetry breaking, requiring additional scrutiny of the nature of this transition. This is work in progress in our group that will be reported in the near future.
@@
Finally, dynamical response to oscillatory or any form of dynamical straining are known to show dipole screening as well, cf. Ref. Hentschel et al. (2024). The presence of Chiral symmetry breaking and it consequences needs to be examined for such protocols as well.
@@ Appendix A Analytic solutions of the equations
Substituting the values of $Z$ from above equations into Eqn.(A), we obtain the values of $\kappa$ defining the transformation in Eqn.(34). Therefore, from equations (32, and 34) we have
@@
Thus Eqns. (66 and 67) reduce to
@@
Now we can write the values of $m_{1}$ and $m_{2}$ from Eqs.(63 and 65)